OpenStudy (anonymous):

Probability is one of the new chapter for me . and I did't get the concept . A bag contains 2 red, 3 green and 2 blue balls. Two balls are drawn at random. What is the probability that none of the balls drawn is blue? A. 10 21 B. 11 21 C. 2 7 D. 5 7

OpenStudy (anonymous):

@mayankdevnani @imqwerty

OpenStudy (anonymous):

Are the balls replaced after being drawn?

OpenStudy (anonymous):

no

OpenStudy (welshfella):

from the choices it would seem they are not replaced

OpenStudy (anonymous):

Okay, so let's start with the first ball. The bag currently has 7 balls in it, of which 2 are blue, so the probability of drawing a blue ball is \(\frac{2}{7}\). Do you follow so far?

OpenStudy (anonymous):

yeah! got it

OpenStudy (anonymous):

Okay, now can you tell me the probability of the second ball taken being blue? Remember that the bag is now missing one blue ball.

OpenStudy (anonymous):

1/6 ?

OpenStudy (anonymous):

Exactly, good work. Now we have the probability of two events happening, do you know how to work out the probability of both happening?

OpenStudy (anonymous):

no idea :(

OpenStudy (anonymous):

Oh crap, completely misread your question. Ignore that. Let's wind back a bit here.

OpenStudy (anonymous):

:O ok

OpenStudy (anonymous):

I read it as both blue when it's neither blue. Okay, well now you have a little understanding about probability, can you tell me the probability of the first ball NOT being blue?

OpenStudy (anonymous):

wild guess !! 5/7 ?

OpenStudy (anonymous):

Exactly, well done. There are 5 balls in the bag that aren't blue, out of the 7. What about the second ball?

OpenStudy (anonymous):

4/6 ?

OpenStudy (anonymous):

Yep, after removing one of the non-blue balls we're left with 6 balls in total, 4 of which aren't blue. Can you read this and see if you can tell me the probability of both happening: http://www.bbc.co.uk/schools/gcsebitesize/maths/statistics/probabilityhirev2.shtml

OpenStudy (anonymous):

i think it should be , 1/7 * 1/6 = 1/42

OpenStudy (anonymous):

You're correct about the multiplication, P(A and B)=P(A)P(B) but your answer is wrong -this is partly my fault for me confusing you. You've used the wrong values, it should be (5/7)*(4/6).

OpenStudy (anonymous):

:( did't get this

OpenStudy (anonymous):

1/7 and 1/6 were the probabilities we worked out before for drawing a blue ball then a blue ball - this wasn't the right thing to work out and is my fault for misreading the question. We then worked out the probability of the first ball being NOT blue as 5/7, and then we worked out the probability of the second ball being NOT blue as 4/6. So the probability of these two events both happening is (5/7)*(4/6). Do you understand now?

OpenStudy (anonymous):

oh ! yeah got it .

OpenStudy (anonymous):

thanks for being patient and teach me @tom982 :)

OpenStudy (welshfella):

another way of doing this wouild be to draw a tree diagran. Have you faniliar with them?

OpenStudy (anonymous):

nope

OpenStudy (welshfella):

you probably be introduced to them sometime

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

No problem. I was tempted to do the tree diagram @welshfella but wasn't sure if it had been covered yet.

OpenStudy (welshfella):